8 arn ey ya nyt ya5 25 up se li Me lah ll op of oodi ? tbh m nm
Answer #1

1) Option A is correct. This can be explained by the table below:

Age Years lived Income Income-Expenditure

5 35    0+800,000+0=$800,000 800,000-700,000 =100,000

15 25   0+600,000=$600,000 600,000-500,000=100,000

25 15   0+200,000=$200,000 200,000-300,000= -100,000

35 5 0 0-100,000= -100,000

Thus, we can say that annual distribution of net income is equal to the line of equality.

2) Option D is correct as total value of claims = $100,000*10 =$1,000,000 = $1 million